OpenStudy (horrorfreakk):

Can someone help me with this question? Meg has a can that contains 80% of orange juice and the rest water. The can has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of orange juice and water in the can. Define the variable used in the equation. Part B: How many liters of orange juice are present in the can? Show your work.

OpenStudy (dumbcow):

80% +20% = 100% Let x be total amount in liters \[0.8x + 0.2x = x\] \[0.8x + 1 = x\] solve for x

OpenStudy (horrorfreakk):

0.8 + 0.2 = 1 and 0.8 + 1 = 1.8